the area of a rectangle wall of a barn is 171 square feet. Its length is 10 feet longer than its width. Find the length and width of the wall of the barn.

Answer by jorel1380(3719)

w(w+10)=171
w2+10w=171
w2+10w-171=0
(w+19)(w-9)=0
w=-19 or 9
w+10=19
Throwing out the negative answer, we get the dimensions of the barn wall to be 19' by 9'. Try and hit that..